Solve.
The square root of x-1 = x-3
Can someone please explain this? Thank you.

Answer by checkley75(3666)

SQRT(X-1)=X-3 NOW WE SQUARE BOTH SIDES OF THIS EQUATION.
X-1=X^2-6X+9
X^2-6X-X+9+1=0
X^2-7X+10=0
(X-5)(X-2)=0
X-5=0
X=5 ANSWER.
X-2=0
X=2 ANSWER (but see below!).
PROOFS
SQRT(5-1)=5-3
SQRT4=2
2=2
SQRT(2-1)=2-3
sqrt1=-1
1=-1 THIS ANSWER MUST BE REJECTED!!